# On the kinks, the solitons and the shocks in discrete nonlinear transmission line

### 24 Feb 2024 v5

@@ Abstract
We studied lossless discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors. The travelling waves with finite support in such systems are the kinks and the solitons, which speeds and profiles were calculated. We also studied lossy transmission lines. The travelling waves with finite support in such systems are the dissipative kinks and the shocks, which speeds and profiles were also calculated.
@@ II.1 The travelling waves
Figure 1: Discrete transmission line.
@@
We’ll consider travelling waves with finite support, that is impose upon $V(Z)$ the boundary conditions
@@ III The dissipative kinks and shocks
Let us ask ourselves what happens with the kinks when the transmission line becomes lossy? To answer this question let us include in the system the resistors $R$ shunting the inductors (see Fig. 2).
Figure 2: Discrete transmission line with shunted inductors.
@@
The point $V=V_{1}$ should be the position of unstable equilibrium. The point $V=V_{2}$ can be the point of either stable, or unstable equilibrium. (We met with similar dichotomy while studying travelling waves with finite support in Josephson transmission line kogan4 ). The waves corresponding to the stable equilibrium at the point $V=V_{2}$ we called the shocks, to the unstable equilibrium – the kinks.
@@ IV Conclusions
We considered finite support travelling waves of the voltage and the current in the discrete lossless transmission line with nonlinear capacitors. We have shown that these waves are of two kinds: the kinks, when the voltage has different asymptotic values at both sides of the transmission line, and the solitons (solitary waves), when the voltage has the same asymptotic value at both sides of the transmission line. We also studied lossy transmission lines. The travelling waves with finite support in such systems are the dissipative kinks and the shocks. We calculated the speeds and the profiles both of the dissipative kinks and of the shocks.

### 19 Mar 2024 v6

@@ Abstract
We studied discrete transmission lines constructed from ideal linear inductors and nonlinear capacitors (and possibly resistors). The localised travelling waves in the lossless transmission lines are the kinks and the solitons, which speeds and profiles were calculated. The localised travelling waves in the lossy transmission lines are the dissipative kinks and the shocks, which speeds and profiles were also calculated.
@@ II.1 The travelling waves
Figure 1: The lossless discrete transmission line.
@@
We’ll consider localised travelling waves, that is impose upon $V(Z)$ the boundary conditions
@@ III The dissipative kinks and shocks
Let us ask ourselves what happens with the kinks when the transmission line becomes lossy? To answer this question let us include in the system the resistors $R_{L}$ shunting the inductors and the resistors $R_{C}$ in series with the nonlinear capacitors (see Fig. 2).
Figure 2: The lossy discrete transmission line.
@@
The point $V=V_{1}$ should be the position of unstable equilibrium. The point $V=V_{2}$ can be the point of either stable, or unstable equilibrium. (We met with similar dichotomy while studying localised travelling waves in Josephson transmission line kogan4 ). The waves corresponding to the stable equilibrium at the point $V=V_{2}$ we called the shocks, to the unstable equilibrium – the kinks.
@@ IV Conclusions
We considered localised travelling waves of the voltage and the current in the discrete lossless transmission line with nonlinear capacitors. We have shown that these waves are of two kinds: the kinks, when the voltage has different asymptotic values at both sides of the transmission line, and the solitons (solitary waves), when the voltage has the same asymptotic value at both sides of the transmission line. We also studied lossy transmission lines. The localised travelling waves in such systems are the dissipative kinks and the shocks. We calculated the speeds and the profiles both of the dissipative kinks and of the shocks.